OpenStudy (anonymous):

This line is the graph of which equation? (0,-1) (2,0) Answer: a. 2x - y = -1 b. 2x + 2y = 1 c. 2x + y = 1 d. x + 2y = -1 e. x - 2y = 2 i can get y=2x-1 and -2=2(-1/2)-1 but i cannot figure out how to put it in the form the possible answers are in.

OpenStudy (anonymous):

e. x-2y=2

OpenStudy (mathlegend):

@kristin848 can you show your work please? I would like to know how you got that.

OpenStudy (anonymous):

@kristin848 how did you get that answer?

OpenStudy (anonymous):

oh i see , x - 2y = 2 subtract x from both sides and you get: -2y = 2 - x divide both sides by -2 and you get: y = -1 + 2x or y = 2x - 1 , and the equation of the line is y = 2x -1 it matches so e is right , thanks!

OpenStudy (anonymous):

First you need to find the slope, which is (y2-y1)/(x2-x1) (0-(-1))/(2-0)=1/2 then use the point slope formula: y-y1=m(x-x1) y-(-1)=(1/2)(x-0) y+1=(1/2)(x) (get x alone) 2(y+1)=x 2y+2=x 2=x-2y

OpenStudy (anonymous):

the original problem has a picture of the graph it wouldn't let me copy and paste it , so by looking at the graph it was easy to find the slope and the y intercept"b" then use the equation y = mx + b and use a checking method with all the possible answers. thanks!